# SurgLaVi: Large-Scale Hierarchical Dataset for Surgical Vision-Language Representation Learning

### v1

@@ Abstract
Abstract: Vision–language pre-training (VLP) offers unique advantages for surgery by aligning language with surgical videos, enabling workflow understanding and transfer across tasks without relying on expert-labeled datasets. However, progress in surgical VLP remains constrained by the limited scale, procedural diversity, semantic quality, and hierarchical structure of existing datasets. In this work, we present SurgLaVi, the largest and most diverse surgical vision–language dataset to date, comprising nearly 240k clip–caption pairs from more than 200 procedures, and comprising hierarchical levels at phase-, step-, and task-level.
At the core of SurgLaVi lies a fully automated pipeline that systematically generates fine-grained transcriptions of surgical videos and segments them into coherent procedural units. To ensure high-quality annotations, it applies dual-modality filtering to remove irrelevant and noisy samples. Within this framework, the resulting captions are enriched with contextual detail, producing annotations that are both semantically rich and easy to interpret. To ensure accessibility, we release SurgLaVi-, an open-source derivative of 113k clip–caption pairs constructed entirely from public data, which is over four times larger than existing surgical VLP datasets. To demonstrate the value of SurgLaVi datasets, we introduce SurgCLIP, a CLIP-style video–text contrastive framework with dual encoders, as a representative base model. SurgCLIP achieves consistent improvements across phase, step, action, and tool recognition, surpassing prior state-of-the-art methods, often by large margins. These results validate that large-scale, semantically rich, and hierarchically structured datasets directly translate into stronger and more generalizable representations, establishing SurgLaVi as a key resource for developing surgical foundation models.
Keywords: Surgical vision–language pre-training, vision–language models (VLMs), hierarchical video–language datasets, weakly supervised learning, transfer learning, SurgLaVi, foundation models
@@ 1 Introduction
Vision–language pre-training (VLP) aligns visual and textual modalities in a joint embedding space, grounding visual features in language and enabling a wide range of downstream tasks such as open-vocabulary recognition and retrieval [1], captioning [2], visual question answering [3], and spatial grounding [4]. Seminal works like CLIP [1] and SigLIP [5] established this paradigm by training on large-scale image–text pairs, demonstrating strong transfer without task-specific supervision and reducing the dependence on manually labeled datasets. Beyond natural images, this approach is especially valuable in surgery: aligning procedural language with surgical videos facilitates workflow understanding, supports transfer across tasks such as phase, step, and instrument recognition, and mitigates the existing bottleneck of scarce expert-labeled datasets in surgery.
@@
Recent efforts have begun adapting VLP to the surgical domain by segmenting academic surgical videos into clips and pairing them with automatic audio transcriptions [6, 7], demonstrating stronger generalization than models pre-trained exclusively on natural images. Despite these advances, surgical video–language pre-trained models still lag significantly behind task-specific supervised approaches in downstream performance. While increasingly elaborate pre-training frameworks have been proposed to address this gap [7, 8], we argue that the fundamental limitation lies not majorly in model design but in the datasets themselves, which face constraints across several dimensions. First, current pipelines to build VLP datasets rely on heuristic rules to segment academic surgical videos into clips, which frequently truncate or fragment coherent surgical concepts, resulting in clip–caption pairs that lack complete semantic meaning. Although some works have sought to enrich semantics with broader supervision, such as phase-level annotations or video-level abstracts [9, 8], these labels are too coarse to capture the fine-grained, hierarchical structure of surgical workflows, which are composed of actions, tasks, steps, and phases [10]. Second, existing datasets contain a substantial proportion of noisy pairs, meaning the text does not accurately describe the corresponding visual content or contains information irrelevant to surgical workflows. Such noise typically results from temporal misalignment between video and narration due to inaccurate transcription, the inclusion of non-surgical segments present in academic presentations, and narrator commentary that diverges from the surgical scene. Third, existing datasets struggle with scale and diversity. For instance, SVL [6] offers 25k clip–caption pairs but focuses exclusively on laparoscopic procedures, while GenSurgery+ [7] broadens coverage to robotic-assisted surgery, yet remains smaller in scale with only 17k pairs. Finally, none of the existing datasets are openly available, further constraining reproducibility and progress in the field. Together, these issues reveal that advancing surgical VLP requires rethinking not only model pre-training frameworks but, more fundamentally, the construction of datasets that better reflect the semantics, structure, and diversity of surgical workflows.
@@
In response to these limitations, we propose a three-part contribution consisting of: (i) the collection of a large and diverse corpus of surgical videos, and (ii) the design of a novel, fully automated, and scalable pipeline to generate high-quality clip–caption pairs that are temporally precise, semantically rich across different levels of granularity, and representative of real surgical workflows, and (iii) the design of a simple lightweight base model with rigorous and exhaustive experimental study on the dataset, validating that well-curated data enables minimal architectures to surpass more complex models trained on suboptimally designed datasets, highlighting the decisive role of our dataset quality and scale over model complexity. Our dataset pipeline systematically integrates four key components: (i) fine-grained transcription of surgical videos, (ii) semantic hierarchical segmentation into clips such that each pair reflects a complete procedural unit at phase-, step-, and task-level granularity, (iii) dual-modality filtering to discard non-surgical visual content and incomplete or non-descriptive captions, and (iv) contextual enrichment that improves caption quality using preceding narrations and video metadata. Leveraging this pipeline, we build the SurgLaVi dataset, the largest and most diverse surgical vision–language data resource to date, comprising nearly 240k clip–caption pairs from over 200 distinct procedure types, as illustrated in Figure 1. By structuring data hierarchically, SurgLaVi captures complete procedural semantics across multiple temporal scales, enabling representations that are more discriminative at different levels of surgical recognition. Using our automated pipeline, we also construct SurgLaVi, an open-source derivative of 113k clip–caption pairs built entirely from public data, which is more than 4x larger than existing surgical vision–language datasets and is publicly released 222https://github.com/aperezr20/SurgLaVi to foster research.
@@
To assess SurgLaVi’s utility for representation learning, we present SurgCLIP, a lightweight CLIP-style contrastive framework with dual video and text encoders. We pretrain SurgCLIP on SurgLaVi and explore its effectiveness on four downstream tasks: phase, step, action, and tool recognition, spanning seven public laparoscopic and robotic surgical datasets. Evaluation under zero-shot as well as few- and full-shot linear probing demonstrates superior performance. Remarkably, even when pre-trained on SurgLaVi, a smaller subset of the dataset, SurgCLIP surpasses previous state-of-the-art methods across multiple benchmarks. Pretraining on the full SurgLaVi corpus further boosts performance across tasks and modalities, yielding substantial zero-shot F1-score improvements (+10% on Cholec80, +28% on AutoLaparo, and +25% on GraSP) and establishing a new state of the art across all seven datasets. These results highlight that pretraining on SurgLaVi’s hierarchically structured dataset enables a standard model to effectively encode procedural workflows without relying on sophisticated architectures or specialized hierarchical pretraining techniques. Extensive ablation studies further validate the design choices and confirm the robustness of our approach.
@@ 2.1 Surgical Vision-Language Pre-Training Datasets
The scarcity of expert surgical annotations has motivated weakly supervised approaches that pair surgical videos with narrations as contextual grounding. The SVL dataset [6] pioneered this direction, introducing 25k clip–caption pairs from lecture videos using dual automatic speech recognition (ASR) systems: AWS Medical Transcribe [11] and Whisper [12]. While this strategy combined strengths in medical terminology recognition and sentence coherence, it suffered from inaccurate timestamps, leading to temporal misalignment. SVL dataset was further limited by sparse random clip sampling, narrowed diversity by just covering laparoscopic procedures, and the absence of workflow structure. To address this lack of workflow representation, HecVL [9] augmented SVL with phase-level summaries and video-level abstracts, and PeskaVLP [8] leveraged large language models to correct narration errors, expand phase descriptions, and generate abstracts. These extensions enriched textual context but ultimately inherited SVL’s small scale, noisy alignments, and narrowed procedural range. To increase coverage, GenSurg+ [7] turned to publicly available YouTube data, producing 17k clip–caption pairs across 28 robotic and laparoscopic procedures. While this increased procedural diversity, the dataset was even smaller than SVL and relied on a fixed segmentation of 45-second clips, limiting semantic coherence as well. In parallel, ophthalmological surgery demonstrated the potential of large-scale curation with OphVL [13], which assembled 375k pairs through ASR denoising with SurgicBERTa [14] validation and structured metadata from silent videos. OphVL showed that scale and hierarchical structure can drive strong generalization, but its scope remained confined to a single specialty. Overall, these efforts illustrate an iterative progression, each addressing shortcomings of its predecessors, yet surgical VLP datasets remain fundamentally constrained by scale, temporal alignment, procedural diversity, and semantic quality.
@@ 2.2 Vision-Language Pre-training Models in Surgery
Computer vision frameworks such as CLIP [1], ALIGN [15], and SigLIP [5] achieve strong performance through standard contrastive pre-training on large-scale image-text corpora, typically containing millions of pairs. In contrast, the surgical domain has needed creative strategies to achieve representation quality with small, noisy pre-training resources. The foundational SurgVLP [6] demonstrated that domain-specific pre-training outperformed generic CLIP models by learning joint video-text embeddings through dual-branch contrastive learning, yet it could not capture hierarchical workflow structure. To address this limitation, HecVL [9] introduced multi-level contrastive learning across clip-, phase-, and video-level embeddings, creating separate embedding spaces and employing alternating training strategies to capture both short-term actions and long-term procedural context without catastrophic forgetting. However, HecVL’s multi-space approach increased model complexity while struggling with textual quality issues. PeskaVLP [8] tackled the caption quality problem through hierarchical knowledge augmentation using large language models to correct, expand, and summarize texts at different hierarchical levels, while introducing procedure-aware training with Dynamic Time Warping-based contrastive regularization to enforce temporal alignment between video frames and text sequences. Despite these semantic improvements, the approach required complex alternating training schedules and is computationally intensive. Following these approaches and moving towards unified pre-training frameworks, VidLPRO [7] employed a multi-modal fusion approach with cross-attention mechanisms, combining video-text contrastive learning, video-text matching, and masked language modeling objectives in a single stage, though this increased architectural complexity compared to dual-encoder designs. Meanwhile, in ophthalmology, OphCLIP [13] demonstrated retrieval-augmented learning by incorporating silent videos through a memory bank system that matches narrative videos with relevant procedures, achieving hierarchical representation learning through alternating clip- and video-level pretraining. These collective efforts represent increasingly sophisticated architectural approaches to extract maximum representational value from limited supervision. However, model complexity alone cannot overcome fundamental training constraints, with each innovation introducing new computational or methodological trade-offs.
@@ 3 SurgLaVi Dataset and Processing Pipeline
As illustrated in Figure 2, we introduce a novel, fully automated and scalable four-stage pipeline for generating hierarchical video-language datasets from surgical videos. This process leverages different foundation models to perform key operations, including transcription, temporal segmentation, and filtering; thereby generating hierarchical annotations that describe surgical workflows under a language supervision paradigm. The following subsections describe each stage of our pipeline:
@@ 3.1 Stage 1: Speech-to-text conversion
We extracted audio from each video and transcribed it using WhisperX [16], denoted as , a model that generates sentence-level timestamped transcriptions and then adds a forced phoneme alignment to generate millisecond-accurate word-level timestamps. This fine-grained alignment enables flexible and accurate video-text pairing in subsequent stages. Formally, given a video dataset , where is the -th video and is its corresponding audio track, we apply WhisperX to each audio to obtain two terms.
@@ 3.2 Stage 2: Hierarchical Semantic Clip–Caption Pair Generation
Hierarchical semantic segmentation: We employ the GPT-4o large language model (LLM) to restructure sentence-level transcriptions from Stage 1 into semantically coherent segments at three hierarchical levels of surgical granularity, following the SAGES consensus recommendations for temporal event annotation in surgical video [10]. We adopt the phase, step, and task levels from the SAGES framework to represent progressively finer procedural detail, excluding the action level given that atomic actions are difficult to infer from audio alone. For each video’s transcription set , we prompt GPT-4o once (see Fig. 8 in Supplementary Material (SM)) to merge temporally adjacent sentences according to semantic coherence, resulting in hierarchical segmentations:
@@
denotes the phase-level segments (coarse procedural sections),
@@
denotes the step-level segments (mid-level procedural goals),
@@
denotes the task-level segments (fine-grained sets of surgical actions).
@@
This design offers two main benefits. First, it generates video–text pairs at multiple temporal scales: from broad workflow phases to individual surgical tasks, capturing progressively finer procedural detail. Second, it naturally augments the dataset by generating multiple segments from the same video, substantially increasing the diversity and quantity of clip–caption pairs without additional data collection.
@@
We adopt a fine-grained filtering strategy that operates at the clip level, preserving valuable surgical segments rather than discarding entire videos through coarse filtering. This enables preservation of informative surgical segments even when other parts of the same video are irrelevant or noisy. Our goal is to filter out clip–caption pairs that are (i) visually non-surgical or (ii) textually non-descriptive, applying visual and textual checks before retaining a pair.
Visual filtering: We apply visual filtering at the most granular level, and propagate to corresponding step- and phase- level clips to maintain computational efficiency while ensuring hierarchical consistency. For each task-level clip , we uniformly sample frames and classify them using a zero-shot SigLIP vision-language model [17]. We use design prompt templates (Fig. 1 in SM) to define the surgical and non-surgical classes. For each class, we compute its embedding as the average of multiple prompts, which we found empirically to produce more robust predictions than representing each class with a single prompt. Each frame is assigned to the class whose averaged prompt embedding has the highest cosine similarity with the frame embedding.
A task-level clip is classified as surgical if more than of its sampled frames are labeled as surgical:
Once all clips in are classified, labels are propagated directly to their corresponding clips in and to ensure hierarchical consistency. A higher-level clip inherits its label if all of its constituent task segments share the same label. If labels are mixed, the higher-level clip is assigned the majority label among its task segments. This direct propagation ensures consistent filtering throughout the hierarchy while preserving the fine-grained classification performed at the most granular level.
Text descriptiveness filtering: For textual filtering, each caption is processed by , using a prompt (Fig. 9 in SM) to instruct the model to determine whether the caption is describing a visual scene. The output is a binary label: descriptive or non-descriptive.

### v2

@@ Abstract
Abstract: Vision–language pre-training (VLP) offers unique advantages for surgery by aligning language with surgical videos, enabling workflow understanding and transfer across tasks without relying on expert-labeled datasets. However, progress in surgical VLP remains constrained by the limited scale, procedural diversity, semantic quality, and hierarchical structure of existing datasets. In this work, we present SurgLaVi, the largest and most diverse surgical vision–language dataset to date, comprising nearly 240k clip–caption pairs from more than 200 procedures, and featuring hierarchical levels at coarse-, mid-, and fine-level. At the core of SurgLaVi lies a fully automated pipeline that systematically generates fine-grained transcriptions of surgical videos and segments them into coherent procedural units. To ensure high-quality annotations, it applies dual-modality filtering to remove irrelevant and noisy samples. Within this framework, the resulting captions are enriched with contextual detail, producing annotations that are both semantically rich and easy to interpret. To ensure accessibility, we release SurgLaVi-, an open-source derivative of 113k clip–caption pairs constructed entirely from public data, which is over four times larger than existing surgical VLP datasets. To demonstrate the value of the SurgLaVi datasets, we introduce SurgCLIP, a CLIP-style video–text contrastive framework with dual encoders, as a representative base model. SurgCLIP achieves consistent improvements across phase, step, action, and tool recognition, surpassing prior state-of-the-art methods, often by large margins. These results validate that large-scale, semantically rich, and hierarchically structured datasets directly translate into stronger and more generalizable representations, establishing SurgLaVi as a key resource for developing surgical foundation models.
@@ 1 Introduction
Vision–language pre-training (VLP) aligns visual and textual modalities in a joint embedding space, grounding visual features in language and enabling a wide range of downstream tasks such as open-vocabulary recognition and retrieval [clip], captioning [li2022blip], visual question answering [lu2019vilbert], and spatial grounding [gu2021open]. Seminal works like CLIP [clip] and SigLIP [zhai2023sigmoid] established this paradigm by training on large-scale image–text pairs, demonstrating strong transfer without task-specific supervision and reducing the dependence on manually labeled datasets. Beyond natural images, this approach is especially valuable in surgery: aligning procedural language with surgical videos facilitates workflow understanding, supports transfer across tasks such as phase, step, and instrument recognition, and mitigates the existing bottleneck of scarce expert-labeled datasets in surgery.
@@
Recent efforts have begun adapting VLP to the surgical domain by segmenting academic surgical videos into clips and pairing them with automatic audio transcriptions [surgvlp, vidlpro], demonstrating stronger generalization than models pre-trained exclusively on natural images. Despite these advances, surgical video–language pre-trained models still lag significantly behind task-specific supervised approaches in downstream performance. While increasingly elaborate pre-training frameworks have been proposed to address this gap [vidlpro, peskavlp], we argue that the fundamental limitation lies not primarly in model design but in the datasets themselves, which face constraints across several dimensions. First, current pipelines to build VLP datasets rely on heuristic rules to segment academic surgical videos into clips, which frequently truncate or fragment coherent surgical concepts, resulting in clip–caption pairs that lack complete semantic meaning. Although some works have sought to enrich semantics with broader supervision, such as phase-level annotations or video-level abstracts [hecvl, peskavlp], these labels are too coarse to capture the fine-grained, hierarchical structure of surgical workflows, which are composed of actions, tasks, steps, and phases [sages_consensus]. Second, existing datasets contain a substantial proportion of noisy pairs, meaning the text does not accurately describe the corresponding visual content or contains information irrelevant to surgical workflows. Such noise typically results from temporal misalignment between video and narration due to inaccurate transcription tools, the inclusion of non-surgical segments present in academic presentations, and narrator commentary that diverges from the surgical scene. Third, existing datasets struggle with scale and diversity. For instance, SVL [surgvlp] offers 25k clip–caption pairs but focuses exclusively on laparoscopic procedures, while GenSurgery+ [vidlpro] broadens coverage to robotic-assisted surgery, yet remains smaller in scale with only 17k pairs. Finally, none of the existing datasets are openly available, further constraining reproducibility and progress in the field. Together, these issues reveal that advancing surgical VLP requires rethinking not only model pre-training frameworks but, more fundamentally, the construction of datasets that better reflect the semantics, structure, and diversity of surgical workflows.
@@
In response to these limitations, we propose a three-part contribution consisting of: (i) the collection of a large and diverse corpus of surgical videos, and (ii) the design of a novel, fully automated, and scalable pipeline to generate high-quality clip–caption pairs that are temporally precise, semantically rich across different levels of granularity, and representative of real surgical workflows, and (iii) the design of a simple lightweight base model with rigorous and exhaustive experimental study on the dataset, validating that well-curated data enables minimal architectures to surpass more complex models trained on suboptimally designed datasets, highlighting the decisive role of our dataset quality and scale over model complexity. Our dataset pipeline systematically integrates four key components: (i) fine-grained transcription of surgical videos, (ii) semantic hierarchical segmentation that partitions videos into clips representing procedural units at coarse-, mid-, and fine- level granularity, (iii) dual-modality filtering to discard non-surgical visual content and incomplete or non-descriptive captions, and (iv) contextual enrichment that improves caption quality using preceding narrations and video metadata. Leveraging this pipeline, we build the SurgLaVi dataset, the largest and most diverse surgical vision–language data resource to date, comprising nearly 240k clip–caption pairs from over 200 distinct procedure types, as illustrated in Fig. 1. By structuring data hierarchically, SurgLaVi captures complete procedural semantics across multiple temporal scales, enabling representations that are more discriminative at different levels of surgical recognition. Using our automated pipeline, we also construct SurgLaVi, an open-source derivative of 113k clip–caption pairs built entirely from public data, which is more than 4x larger than existing surgical vision–language datasets and is publicly released 222https://github.com/aperezr20/SurgLaVi to foster research.
@@
To assess SurgLaVi’s utility for representation learning, we present SurgCLIP, a lightweight CLIP-style contrastive framework with dual video and text encoders. We pretrain SurgCLIP on SurgLaVi and explore its effectiveness on four downstream tasks: phase, step, action, and tool recognition, spanning seven public laparoscopic and robotic surgical datasets. Evaluation under zero-shot as well as few- and full-shot linear probing demonstrates superior performance. Remarkably, even when pre-trained on SurgLaVi, a smaller subset of the dataset, SurgCLIP consistently exceeds previous state-of-the-art F1 performance on several phase recognition benchmarks, including Cholec80 (+0.3%), AutoLaparo (+17.54%), StrasBypass70 (+6.41%), Heichole (+3.62%) and GraSP (+18.12%). When trained on the full SurgLaVi corpus, the model achieves even stronger results, showing substantial zero shot F1 gains across different benchmarks (+10% on Cholec80, +28% on AutoLaparo, +11.14% on StrasBypass70, +1.1% on BernBypass70, +10.51% on Heichole, and +25% on GraSP). Beyond phase recognition, SurgCLIP also advances performance on step, action, and tool recognition benchmarks. These results highlight that pre-training on SurgLaVi’s hierarchically structured dataset enables a standard model to effectively encode procedural workflows without relying on sophisticated architectures or specialized hierarchical pre-training techniques. Extensive ablation studies further validate the design choices and confirm the robustness of our approach.
@@ 2.1 Surgical Vision-Language Pre-Training Datasets
The scarcity of expert surgical annotations has motivated weakly supervised approaches that pair surgical videos with narrations as contextual grounding. The SVL dataset [surgvlp] pioneered this direction, introducing 25k clip–caption pairs from lecture videos using dual automatic speech recognition (ASR) systems: AWS Medical Transcribe [medicalaws] and Whisper [radford2022whisper]. While this strategy combined strengths in medical terminology recognition and sentence coherence, it suffered from inaccurate timestamps, leading to temporal misalignment. SVL dataset was further limited by sparse random clip sampling, narrowed diversity by just covering laparoscopic procedures, and the absence of workflow structure. To address this lack of workflow representation, HecVL [hecvl] augmented SVL with phase-level summaries and video-level abstracts, and PeskaVLP [peskavlp] leveraged large language models to correct narration errors, expand phase descriptions, and generate abstracts. These extensions enriched textual context but ultimately inherited SVL’s small scale, noisy alignments, and narrowed procedural range. To increase coverage, GenSurg+ [vidlpro] turned to publicly available YouTube data, producing 17k clip–caption pairs across 28 robotic and laparoscopic procedures. While this increased procedural diversity, the dataset was even smaller than SVL and relied on a fixed segmentation of 45-second clips, limiting semantic coherence as well. In parallel, ophthalmological surgery demonstrated the potential of large-scale curation with OphVL [ophclip], which assembled 375k pairs through ASR denoising with SurgicBERTa [SurgicBERTa] validation and structured metadata from silent videos. OphVL showed that scale and hierarchical structure can drive strong generalization, but its scope remained confined to a single specialty. Overall, these efforts illustrate an iterative progression, each addressing shortcomings of its predecessors, yet surgical VLP datasets remain fundamentally constrained by scale, temporal alignment, procedural diversity, and semantic quality.
@@ 2.2 Vision-Language Pre-training Models in Surgery
Computer vision frameworks such as CLIP [clip], ALIGN [jia2021scaling], and SigLIP [zhai2023sigmoid] achieve strong performance through standard contrastive pre-training on large-scale image-text corpora, typically containing millions of pairs. In contrast, the surgical domain has needed creative strategies to achieve representation quality with small, noisy pre-training resources. The foundational SurgVLP [surgvlp] demonstrated that domain-specific pre-training outperformed generic CLIP models by learning joint video-text embeddings through dual-branch contrastive learning, yet it could not capture hierarchical workflow structure. To address this limitation, HecVL [hecvl] introduced multi-level contrastive learning across clip-, phase-, and video-level embeddings, creating separate embedding spaces and employing alternating training strategies to capture both short-term actions and long-term procedural context without catastrophic forgetting. However, HecVL’s multi-space approach increased model complexity while struggling with textual quality issues. PeskaVLP [peskavlp] tackled the caption quality problem through hierarchical knowledge augmentation using large language models to correct, expand, and summarize texts at different hierarchical levels, while introducing procedure-aware training with Dynamic Time Warping-based contrastive regularization to enforce temporal alignment between video frames and text sequences. Despite these semantic improvements, the approach required complex alternating training schedules and is computationally intensive. Following these approaches and moving towards unified pre-training frameworks, VidLPRO [vidlpro] employed a multi-modal fusion approach with cross-attention mechanisms, combining video-text contrastive learning, video-text matching, and masked language modeling objectives in a single stage, though this increased architectural complexity compared to dual-encoder designs. Meanwhile, in ophthalmology, OphCLIP [ophclip] demonstrated retrieval-augmented learning by incorporating silent videos through a memory bank system that matches narrative videos with relevant procedures, achieving hierarchical representation learning through alternating clip- and video-level pretraining. These collective efforts represent increasingly sophisticated architectural approaches to extract maximum representational value from limited supervision. However, model complexity alone cannot overcome fundamental training constraints, with each innovation introducing new computational or methodological trade-offs.
@@ 3 SurgLaVi Dataset and Processing Pipeline
As illustrated in Figure 2, we introduce a novel, fully automated and scalable four-stage pipeline for generating hierarchical video-language datasets from surgical videos. This process leverages different foundation models to perform key operations, including transcription, temporal segmentation, and filtering; thereby enabling generation of hierarchical annotations that describe surgical workflows under a language supervision paradigm. The following subsections describe each stage of our proposed pipeline:
@@ 3.1 Stage 1: Speech-to-text conversion
To convert speech narrations from videos to text, we use WhisperX (v3) as it demonstrated superior performance within the surgical domain compared to other ASR models (Table 1 in Appendix). This model, denoted as , generates sentence-level timestamped transcriptions and then adds a forced phoneme alignment to generate millisecond-accurate word-level timestamps. This fine-grained alignment enables flexible and accurate video-text pairing in subsequent stages. Formally, given a video dataset , where is the -th video and is its corresponding audio track, we apply WhisperX to each audio to obtain two terms.
@@ 3.2 Stage 2: Hierarchical Semantic Clip–Caption Pair Generation
Hierarchical semantic segmentation: We employ the GPT-4o large language model (LLM) to restructure sentence-level transcriptions from Stage 1 into semantically coherent segments at three hierarchical levels of surgical granularity. Inspired by the SAGES framework [sages_consensus], which represents procedural activities through progressively finer units of analysis, we adopt a hierarchical approach in which segments reflect increasingly granular procedural goals. In our formulation, these are expressed as coarse, mid, and fine segments, enabling a principled temporal decomposition of the surgical workflow that captures high-level context and fine-grained detail.
For each video’s transcription set , we prompt GPT-4o once (Fig. 8 in Appendix) to merge temporally adjacent sentences according to semantic coherence, resulting in hierarchical segmentations:
@@
denotes the coarse-level segments capturing high-level complex procedural structure,
@@
denotes the mid-level segments reflecting intermediate procedural goals,
@@
denotes the fine-level segments representing localized and more simple surgical tasks.
@@
This design offers two main benefits. First, it generates video–text pairs at multiple temporal scales capturing progressively finer procedural detail. Second, it naturally augments the dataset by generating multiple segments from the same video, substantially increasing the diversity and quantity of clip–caption pairs without additional data collection.
@@
We adopt a fine-grained filtering strategy that operates at the clip level, preserving valuable surgical segments rather than discarding entire videos through coarse filtering. This enables preservation of informative surgical segments even when other parts of the same video are irrelevant or noisy. Our goal is to filter out clip–caption pairs that are (i) visually non-surgical or (ii) textually non-descriptive, applying visual and textual checks before retaining a pair. We describe our proposed filtering methodology for each modality in this section and detailed quantitative results in Sec. 1.1 from Appendix.
Visual filtering: We apply visual filtering at the most granular level, and propagate to corresponding mid- and coarse- level clips to maintain computational efficiency while ensuring hierarchical consistency. For each fine-level clip , we uniformly sample frames and classify them using a zero-shot SigLIP vision-language model [open_clip]. We design prompt templates (Fig. 1 in Appendix) to define the surgical and non-surgical classes. For each class, we compute its embedding as the average of multiple prompts, which we found empirically to produce more robust predictions than representing each class with a single prompt. Each frame is assigned to the class whose averaged prompt embedding has the highest cosine similarity with the frame embedding.
A fine-level clip is classified as surgical if more than of its sampled frames are labeled as surgical:
Once all clips in are classified, labels are propagated directly to their corresponding clips in and to ensure hierarchical consistency. A higher-level clip inherits its label if all of its constituent fine segments share the same label. If labels are mixed, the higher-level clip is assigned the majority label among its fine segments. This direct propagation ensures consistent filtering throughout the hierarchy while preserving the fine-grained classification performed at the most granular level.
Text descriptiveness filtering: For textual filtering, each caption is processed by , using a prompt (Fig. 9 in Appendix) to instruct the model to determine whether the caption is describing a visual scene. The output is a binary label: descriptive or non-descriptive.